OpenStudy (bookworm14):

Teesha is in the French club. There are 31 students in the club. The French teacher will pick two students at random to guide visiting students from France. What is the probability that Teesha will be picked as a guide? Hint: Teesha must be picked first or second. a. 1/30 b. 30/31 c. 1/930 d. 2/31 In my mind I am thinking the first time a name is drawn, there is a 1/31 chance her name will be drawn. the second time there is only a 1/30 chance her name will be picked. I multiplied those and got 1/930 Then again, I am also thinking there are 2/31 chances her name could be drawn.

OpenStudy (welshfella):

what is the probabilty that Teesha will be picked first?

OpenStudy (bookworm14):

1/31

OpenStudy (welshfella):

she is 1 of 31 students yes

OpenStudy (welshfella):

now the other favourable scenario is that she is picked second the probability that she is not picked first time is 30/31 what is the proability she is picked on the second choice?

OpenStudy (welshfella):

- after not being picked on first choice

OpenStudy (bookworm14):

wouldn't it be 1/30 ? because there will only be 30 options left if someone else got picked the first time

OpenStudy (welshfella):

yes

OpenStudy (perfect14):

no it would be 2/30 because she has two chances to be picked

OpenStudy (welshfella):

So probabilty first choice = 1/31 probability chosen second = 30/31 * 1/30 Required probabilty is the sum of these 2.

OpenStudy (welshfella):

No 1/30 was right

OpenStudy (welshfella):

a tree diagram is a good way of doing this one

OpenStudy (bookworm14):

ok, my answers i was torn between when i initially posted this question were C and D, and I posted my thought process on how I came to either of those conclusions. were either of those right?

OpenStudy (welshfella):

P(picked first) = 1/31 P( not picked then picked = 30/31 8 1/30 = 1/31 required probability = 1/31 + 1/31

OpenStudy (welshfella):

* that 8 should be a *

OpenStudy (welshfella):

the 2 scenarios cannot both occur so the probabilities are added

OpenStudy (bookworm14):

ok, that made way more sense, somehow i kept getting 29 based on ur first description ( i was typin it in wrong) so D is the answer?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

when its P1 OR P2 ADD P's when its P1 AND P2 multiply the P's

OpenStudy (bookworm14):

ok thank you!